OpenStudy (s3a):

Why is the following a homogeneous differential equation of degree two?: x sin y/x (y dx + x dy) + cos y/x (x dy - y dx) = 0

OpenStudy (amistre64):

if you can write it in the form such that y/x is a simplified version; then it can be veed out i think

OpenStudy (lgbasallote):

maybe because you have two variables in each monomial? dunno. might be...

OpenStudy (amistre64):

degree 2 would imply a ^2 someplace

OpenStudy (s3a):

degree two means derivative of power 2 in the case of differential equations. I don't see a second order derivative around as for the homogeneity, 2 secs.

OpenStudy (amistre64):

its homo becasue its got a =0 on the end

OpenStudy (s3a):

amistre64, are you sure about the =0 thing you just said?

OpenStudy (amistre64):

they use 2 defs for homogenous in diffy qs; this one conforms to the =0 def

OpenStudy (s3a):

what's the other def?

OpenStudy (amistre64):

f(tx,ty) = t f(x,y)

OpenStudy (s3a):

o that's the order 0 homo eq, right?

OpenStudy (amistre64):

if you attach a random variable such as "t" to all your xs and ys; and can factor it out cleanly; its homogenous

OpenStudy (amistre64):

i got no idea what the name is lol

OpenStudy (s3a):

it's alright, wikipedia said it, i was asking for confirmation

OpenStudy (s3a):

:)

OpenStudy (amistre64):

if wiki says it; it MUST be true lol

OpenStudy (s3a):

lol

OpenStudy (s3a):

it's not as unreliable as people think, in my opinion

OpenStudy (s3a):

there is a (small) team of people constantly watching excluding the rest of the world

OpenStudy (amistre64):

they say a chain is only as strong as its weakest link ... but yeah, ive never found anything glaringly false on it yet

OpenStudy (s3a):

they also say a bodybuilder is as strong as his weakest muscle lol

OpenStudy (s3a):

but ya, why is it degree two?

OpenStudy (amistre64):

can you algebra it into a ^2 on it somplace?

OpenStudy (s3a):

well i can make x^2 but that's not a derivative.

OpenStudy (s3a):

like x^2 vs d^2 x/dt is not the same thing

OpenStudy (s3a):

(random example)

OpenStudy (s3a):

lol, i'll see if i can make it understand

OpenStudy (amistre64):

degree is not order; d^2/dx^2 is an order, not a degree

OpenStudy (s3a):

ooo

OpenStudy (s3a):

but wait then i'd need y^2 rather than x^2, no?

OpenStudy (amistre64):

its not a wiki, but its got an edu at the end ;)

OpenStudy (turingtest):

I was able to get it into the form dx/dy=(something w/ x^2 in it) so I don't see why it has to be y^2

OpenStudy (amistre64):

maybe multiply thru by d/dx

OpenStudy (amistre64):

um, thats just taking the derivative of it tho, maybe lol \[x sin \frac yx (y dx + x dy) + cos \frac yx (x dy - y dx) = 0\] \[x sin \frac yx (y + x \frac{dy}{dx}) + cos \frac yx (x\frac{dy}{dx} - y) = 0\] \[ yx sin \frac yx + xx sin \frac yx \frac{dy}{dx} + xcos \frac yx\frac{dy}{dx} - ycos \frac yx = 0\] \[ yx sin \frac yx + \frac{dy}{dx} (x^2 sin \frac yx + xcos \frac yx) - ycos \frac yx = 0\]

OpenStudy (s3a):

ya d/dx is just taking the derivative ;) and still, all i see is x^2

OpenStudy (s3a):

o i just saw turingtest's answer, well: y is the dependent variable, x is independent. that doesn't matter?

OpenStudy (turingtest):

It says that in the problem?

OpenStudy (s3a):

What? That's it's homogeneous of degree two? Yes, the solution says that.

OpenStudy (turingtest):

no, that y is the dependent variable

OpenStudy (s3a):

No, it doesn't but it seems to reat it as the dependent variable especially with the y = xv substitution.

OpenStudy (amistre64):

are you sure youve notated it correctly?

OpenStudy (s3a):

treat* and what do you mean "notated". do you mean copied it down correctly?

OpenStudy (amistre64):

yes, to notate is to write down

OpenStudy (turingtest):

I'm looking at something that says the degree is the exponent on the derivative, so we seem to be looking for\[(\frac{dy}{dx})^2\]somehow...

OpenStudy (s3a):

and i just double checked and ya, i copied it down correctly?

OpenStudy (s3a):

TuringTest: that's order i think (not degree)

OpenStudy (amistre64):

no, thats degree

OpenStudy (amistre64):

\[\frac{d^2y}{dx^2};order\ 2\] \[\left(\frac{dy}{dx}\right)^2;degree\ 2\]

OpenStudy (turingtest):

yes, I meant the second one there^

OpenStudy (amistre64):

wolfram apparently wants money to work now so its not going to be helpful

OpenStudy (s3a):

no, i think it's just for new features.

OpenStudy (s3a):

there's always open source software like maxima

OpenStudy (amistre64):

new features like; "take your time and solve this" feature lol

OpenStudy (s3a):

lol

OpenStudy (amistre64):

\[yx sin \frac yx + \frac{dy}{dx} (x^2 sin \frac yx + xcos \frac yx) - ycos \frac yx = 0\] \[y=vx;\ \frac{dy}{dx}=\frac{dv}{dx}x+v\] \[vx^2 sin (v) + (\frac{dv}{dx}x+v) (x^2 sin (v) + xcos (v)) - vxcos(v) = 0\] \[vx^2 sin (v) + \frac{dv}{dx}x^3sin(v)+\frac{dv}{dx}x^2cos(v)+vx^2sin(v)+vxcos(v) - vxcos(v) = 0\] \[vx^2 sin (v) + \frac{dv}{dx}(x^3sin(v)+x^2cos(v))+vx^2sin(v)+vxcos(v) - vxcos(v) = 0\] still cant get a deg2

OpenStudy (amistre64):

\[\frac{dv}{dx}(x^3sin(v)+x^2cos(v))+2vx^2sin(v) = 0\] \[\frac{dv}{dx}(x^3sin(v)+x^2cos(v))=-2vx^2sin(v)\] \[\frac{dv}{dx}=-2\frac{vx^2sin(v)}{x^3sin(v)+x^2cos(v)}\] and cant seem to factor out a x^n from the underside to make it homogenous eaither

OpenStudy (s3a):

The book implies it's of second degree prior to the substitution.

OpenStudy (amistre64):

my book says that a dependant set of vectors is independant; so I have to kinda have to wonder how much faith to put into book stuff

OpenStudy (s3a):

so, what does that mean for my problem?

OpenStudy (amistre64):

it means that we cant seem to adapt the problem and conform it to the information given of homogenous and dergree 2. either the solution is : not exist or its simply a bad type up in the book

OpenStudy (amistre64):

ugh, we also got another definition floating out there: f(tx,ty) = t^n f(x,y) is a homogenous eg of degree "n"

OpenStudy (amistre64):

so, attach a t to all your xs and ys and see if we can factor it our cleanly

OpenStudy (amistre64):

\[tx sin ty/tx (ty dx + tx dy) + cos ty/tx (tx dy - ty dx) = 0\] \[tx sin y/x (ty dx + tx dy) + cos y/x (tx dy - ty dx) = 0\] \[tx sin y/x t(y dx + x dy) + cos y/x t(x dy - y dx) = 0\] \[t^2 x sin y/x (y dx + x dy) + tcos y/x (x dy - y dx) = 0\] \[t(t x sin y/x (y dx + x dy) + cos y/x (x dy - y dx) = 0\] we are still stuck with a "t" inside

OpenStudy (amistre64):

you sure you typed it up correctly?

OpenStudy (amistre64):

there should be a: y cos y/x or maybe even x cos y/x to attach another t onto to pull it out clean

OpenStudy (amistre64):

or if that: x sin y/x was just sin y/x but then thats just a t^1 and not a t^2 that gets factored out

OpenStudy (s3a):

Oh, I see. The problem seems well solved so it might be just this part with a typo. Thanks for all this and sorry for exhausting you with "proofs." Yes, I double checked what I typed.

OpenStudy (amistre64):

good luck :)

OpenStudy (s3a):

Thanks again :).